# Optimization tools for distance-preserving flag fault-tolerant error correction

### v3

@@ I Introduction
In a non-ideal FTQEC protocol, faults can occur at any gate on any round of the syndrome measurements. The only information about the error on data qubits that we can obtain is a sequence of error syndromes, and we want to find an appropriate recovery operator from this information. An ideal strategy would be using all syndrome bits from all rounds, that is, the whole measurement outcomes in space-time. For some codes with a nice structure, such as surface codes, an efficient space-time decoder exists [39]. However, constructing a space-time decoder for a general stabilizer code is not simple. To simplify the problem, we will consider an error decoder, which is composed of two parts: the space and the time decoders. Under the assumption that the syndrome measurements can be faulty, the time decoder finds a round of syndrome measurements that has no faults and gives a correct syndrome. The space decoder then uses the correct syndrome to construct a recovery operator.
@@
In this work, we build several optimization tools for both space and time decoders for the purpose of reducing the overhead of both to obtain better-performing protocols for flag qubit-based FTQEC. Most of our tools are applicable to general stabilizer codes, but we primarily focus on a self-dual CSS code in which the number of physical qubits is odd, the number of logical qubits is 1, and logical $X$ and $Z$ operators are transversal for simplicity. Our main results are the following: (1) We develop a technique to build a lookup table more efficiently. Our compact lookup table for a self-dual CSS code requires 87.5% less memory footprint compared to a lookup table for a generic stabilizer code. Our method also efficiently verifies whether a configuration of the flag circuits preserves the code distance. The development also leads to a notion of fault code, which can be useful in error sampling for the circuit-level noise model. (2) We introduce the Meet-in-the-Middle (MIM) technique, which can help the lookup table decoder correct faults more than the number of errors correctable by the underlying code. Although the correction is not always successful, the higher success probability can significantly increase the pseudothreshold in our simulations. (3) We generalize previous work [38] on adaptive syndrome measurement schemes to flag FTQEC and introduce one-tailed and two-tailed adaptive time decoders, which are useful in different circumstances. We also develop a classical processing technique on flag information that makes our FTQEC protocols compatible with any fault-tolerant Clifford computation. (4) We use our optimization tools and perform numerical simulations on the hexagonal color codes [41] of distances 3, 5, 7, and 9. The results show that each of our tools can significantly reduce the logical error rates and increase the pseudothreshold for each code while preserving the code distance. For the hexagonal color code of distance 9, the pseudothreshold is improved by one order of magnitude, from $(1.34\pm 0.01)\times 10^{-4}$ to $(1.42\pm 0.12)\times 10^{-3}$, when all techniques are applied.
@@ III.1 Compact lookup table for minimum weight decoding and fault code
In this section, we discuss how to construct the fault set $\mathcal{F}_{t}$, verify its distinguishability, and construct the lookup table for error decoding. With our method, we can reduce the memory footprint requirement of the lookup table by 87.5% for self-dual CSS codes compared to generic stabilizer codes. We also present the framework of fault codes that enables fast construction using streamlined Pauli-frame simulation represented as matrix algebra operations over $\mathrm{GF}(2)$.
@@ III.1.1 Reducing the memory footprint
Leveraging the structure of CSS codes, we can significantly improve the memory footprint. Assuming standard CSS decoding in which two separate lookup tables are used for $X$ and $Z$ decoding. Denote with $r_{X}$ and $r_{Z}$ the number of $X$- and $Z$-type stabilizer generators satisfying $r_{X}+r_{Z}=n-k$. The per entry cost decreases, as the entries only need to cater for $X$- or $Z$-type operators and syndromes. Each entry for the $X$- and $Z$-type syndromes will have $2r_{X}$ and $2r_{Z}$ bits respectively for the syndrome and the cumulative flag vector, and $n$ bits for the recovery operator. A self-dual CSS code needs only one table futher decreasing the cost (see more detail on the total number of bits in Appendix A). Moreover, we can reduce the number of bits for the recovery operator to $k$ using the following two key ideas:
@@
    Given a syndrome $\vec{s}(E)$, the goal of decoding is to find a recovery operator $R$ such that $RE\in\mathcal{S}$, thus $R$ converts the error into the logical identity operation. For any possible Pauli error, we only have to store its logical class, a value that indicates how the error is related to a CRO with the same syndrome. This enables the map value to be only $2k$ bits of information in general and $k$ bits in case the code is a self-dual CSS code. In this latter case and with $k=1$, the logical class is 0 if the multiplication of the Pauli operator and the CRO with the same syndrome is in the stabilizer group, otherwise the logical class is 1.
Altogether, for a self-dual CSS code with $n\gg k$, the size of the table can be as small as 12.5% of the table if we viewed the code as a generic stabilizer code and stored the full recovery operators instead of the logical classes. For a CSS code that is not self-dual, the gain is smaller but still significant. See Appendix A for detailed calculations around savings for the lookup table. Note that if the lookup table is used for proving distinguishability, all unique syndromes are required. However, in a realtime decoding architecture, the entries corresponding to significantly low probability fault combinations may be excluded, resulting in further reduction [47].
@@ III.1.2 Constructing the lookup table
Enumerating weight-1 faults— From here on, we will only consider a self-dual CSS code, and denote its parity check matrix $H$. In order to list all possible single faults under the circuit-level depolarizing noise model, it is sufficient to consider all possible weight-1 faults within a single round of syndrome measurements. Each column of the fault check matrix $H_{\text{f}}$ describes for each possible weight-1 fault what its full syndrome and its logical class are. As the logical class of each fault depends on how its CRO is defined, we define the fault check matrix relative to the right inverse $H^{-1}$ of $H$ (for which $HH^{-1}=I_{(n-k)/2}$).
@@
Generalizing $H_{\text{f}}$ for a non-self-dual CSS code is straightforward. In that case, the parity check matrices for $X$- and $Z$-type errors can be different, leading to different fault check matrices. Generalizing $H_{\text{f}}$ for a generic stabilizer code is more complicated but still doable, as all operators must be considered in the symplectic form. In that case, the number of rows for the logical class is $2k$. Also, instead of taking the inner product with $J_{i}$, whether each CRO commutes or anticommutes with each logical operator can be determined by the symplectic inner product between the symplectic bitstrings representing the CRO and the logical operator.
In the case that the code is a self-dual CSS code, $n$ is odd, $k=1$, and logical $X$ and logical $Z$ operators are transversal, the fault check matrix is,
${\color[rgb]{0,0,0}H_{\text{f}}}=\left(\begin{array}[]{c|c|c}H&0&\hbox{\multirowsetup$H_{\text{f,gate}}$}\\ 0&I_{(n-1)/2}&\\ J^{T}(I_{n}\oplus H^{-1}H)&0&\end{array}\right),$
@@
For errors of weight 1 on the data qubits, the operator $RE$ of each error can be represented by each column of $I_{n}\oplus H^{-1}H$. Since the logical class of $RE$ can be determined by its weight parity, the logical classes of this type of errors are the row of $L\equiv J^{T}(I_{n}\oplus H^{-1}H)$ where $J$ is the all-one column vector. That is for the Steane code, the part of ${\color[rgb]{0,0,0}H_{\text{f}}}$ corresponding to pure data qubit errors is
@@
$P_{i}=\begin{pmatrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&{\color[rgb]{0,0,0}1}&0&{\color[rgb]{0,0,0}0}\\ 0&0&1&0&0&0\\ 0&0&0&{\color[rgb]{0,0,0}0}&0&{\color[rgb]{0,0,0}1}\\ 0&1&0&0&1&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{pmatrix}.$ (18)
@@
$PA=\begin{pmatrix}{\color[rgb]{0,0,0}\Omega}\\ \hline\cr\Phi\end{pmatrix}=\begin{pmatrix}{\color[rgb]{0,0,0}\Omega_{1}}&{\color[rgb]{0,0,0}\Omega_{2}}&\ldots&{\color[rgb]{0,0,0}\Omega_{r}}\\ \hline\cr\Phi_{1}&\Phi_{2}&\ldots&\Phi_{r}\end{pmatrix},$ (30)
@@
$L=J^{T}({\color[rgb]{0,0,0}\Omega}\oplus\Theta).$ (33)
As a result, the part of ${\color[rgb]{0,0,0}H_{\text{f}}}$ corresponding to the gate faults is,
$H_{\text{f,gate}}=\begin{pmatrix}S\\ \hline\cr\Phi\\ \hline\cr L\end{pmatrix}=\begin{pmatrix}H{\color[rgb]{0,0,0}\Omega}\\ \hline\cr\Phi\\ \hline\cr J^{T}(I_{n}\oplus H^{-1}H){\color[rgb]{0,0,0}\Omega}\end{pmatrix}.$ (40)
The relationship between the full syndrome, the data error, and the CRO of each fault is as follows: Suppose that the $i$-th column of ${\color[rgb]{0,0,0}H_{\text{f}}}$ (which represents a single fault on the $i$-th location) contains error syndrome $\vec{s}_{i}$, flag vector $\vec{f}_{i}$, and logical class $l_{i}$. The CRO of the fault is $H^{-1}\vec{s}_{i}$, while the data error of the fault is $l_{i}J\oplus H^{-1}\vec{s}_{i}$. That is, in case of a single fault, the actual recovery operator (ARO) we need to apply when finding the full syndrome $(\vec{s}_{i},\vec{f}_{i})$ is $l_{i}J\oplus H^{-1}\vec{s}_{i}$.
@@
When $t\geq 2$, we populate the cache with the logical classes of higher-weight fault combinations by simply combining all possible fault combinations of lower-weight fault combinations while keeping track of the weights of the fault combinations. We describe the $i$-th fault combination as a key-value pair $[(\vec{s}_{i},{\color[rgb]{0,0,0}\vec{F}_{i}}):(l_{i},w_{i})]$ where $(\vec{s}_{i},{\color[rgb]{0,0,0}\vec{F}_{i}})$ is the full syndrome, $l_{i}$ is the logical class, and $w_{i}$ is the weight of the fault combination. Combining the $i$-th and the $j$-th fault combinations gives $[(\vec{s}_{i}\oplus\vec{s}_{j},{\color[rgb]{0,0,0}\vec{F}_{i}}\oplus{\color[rgb]{0,0,0}\vec{F}_{j}}):(l_{i}\oplus l_{j},w_{i}+w_{j})]$. As we aim to check whether $\mathcal{F}_{t}$ is distinguishable, we fill up the cache by combining any pair of fault combinations that satisfy $w_{i}+w_{j}\leq t$. In case that the process gives the new key (the full syndrome) that already exists in the cache, we have a key conflict. This can be one of the following cases:
@@
If at the end we find that $\mathcal{F}_{t}$ is distinguishable, we can construct a lookup table of search radius $t$ from the cache as follows: for each key-value pair $[(\vec{s}_{i},{\color[rgb]{0,0,0}\vec{F}_{i}}):(l_{i},w_{i})]$ in the cache, we store a new key-value pair $[(\vec{s}_{i},{\color[rgb]{0,0,0}\vec{F}_{i}}):l_{i}J\oplus H^{-1}\vec{s}_{i}]$ in the lookup table (the weights are not necessary for decoding, though it might be useful for estimating the number of faults that causes the full syndrome). That is, $l_{i}J\oplus H^{-1}\vec{s}_{i}$ is the ARO for the full syndrome $(\vec{s}_{i},{\color[rgb]{0,0,0}\vec{F}_{i}})$. When performing error decoding, the ARO is applied if the full syndrome obtained from measurements is found on the lookup table; otherwise, the CRO ($H^{-1}\vec{s}_{i}$) is applied.
@@ III.1.3 The fault code
The perspective of the fault code can also be useful to extend a technique frequently used for error sampling (in qecsim [54], for example) to the circuit-level noise model beyond the code capacity noise model (memory errors only) and phenomenological noise model (both memory and measurement errors). Here, a randomly generated column vector of Hamming weight $w$ now represents faults on $w$ locations instead of errors on $w$ qubits. Suppose that the vector $\vec{v}$ represents the fault combination and $H_{\text{f}}\vec{v}$ gives the full syndrome $(\vec{s}(\vec{v}),{\color[rgb]{0,0,0}\vec{F}}(\vec{v}))$ and the logical class $l(\vec{v})$. In an error correction simulation, the decoder can predict the recovery operator $\vec{r}$ based on the full syndrome. We will find that the predicted recovery operator causes a logical error if and only if $l(\vec{v})$ and $l(\vec{r})$ differ.
@@ IV Optimization Tools for Time decoding
In this section, we develop several types of time decoders for flag FTQEC, building on the ideas of adaptive decoders for Shor-style error correction [38]. Different time decoders use different fault count estimation procedures. In Section IV.1, we describe a conventional way to perform repeated syndrome measurements for flag FTQEC in terms of difference vectors, which will be useful for the development in latter sections. In Section IV.2, we develop one-tailed and two-tailed adaptive time decoders which utilize flag information in the protocols. One-tailed adaptive decoder is applicable to a larger family of codes, while the two-tailed adaptive decoder is more optimized to self-dual CSS codes but need to be used with an extended technique so that it becomes fully fault tolerant when applying to quantum computation. In Section IV.3, we develop two extended techniques that can further improve the performance of our adaptive time decoders for FTQEC, given that the code being used is a self-dual CSS code.

### v4

@@ I Introduction
During the execution of an FTQEC protocol, faults can occur at any gate on any round of the syndrome measurements. The only information about the error on data qubits that we can obtain is a sequence of error syndromes, and we want to find an appropriate recovery operator from this information. An ideal strategy would be using all syndrome bits from all rounds, that is, the whole measurement outcomes in space-time. For some codes with a nice structure, such as surface codes, an efficient space-time decoder exists [39]. However, constructing a space-time decoder for a general stabilizer code is not simple. To simplify the problem, we will consider an error decoder, which is composed of two parts: the space and the time decoders. Under the assumption that the syndrome measurements can be faulty, the time decoder finds a round of syndrome measurements that has no faults and gives a correct syndrome. The space decoder then uses the correct syndrome to construct a recovery operator.
@@
In this work, we build several optimization tools for both space and time decoders for the purpose of reducing the overhead of both to obtain better-performing protocols for flag qubit-based FTQEC. Most of our tools are applicable to general stabilizer codes, but we primarily focus on self-orthogonal CSS codes (CSS codes in which $X$- and $Z$-type generators are of the same form) in which the number of physical qubits is odd, the number of logical qubits is 1, and logical $X$ and $Z$ operators are transversal for simplicity. Our main results are the following: (1) We develop a technique to build a lookup table more efficiently. Our compact lookup table can leverage the structure of a self-orthogonal CSS code and requires 87.5% less memory footprint compared to a lookup table designed for a generic stabilizer code. Our method also efficiently verifies whether a configuration of the flag circuits preserves the code distance. The development also leads to a notion of fault code, which can be useful in error sampling for the circuit-level noise model. (2) We introduce the Meet-in-the-Middle (MIM) technique, which can help the lookup table decoder correct faults more than the number of errors correctable by the underlying code. Although the correction is not always successful, the higher success probability can significantly increase the pseudothreshold in our simulations. (3) We generalize previous work [38] on adaptive syndrome measurement schemes to flag FTQEC and introduce one-tailed and two-tailed adaptive time decoders, which are useful in different circumstances. We also develop a classical processing technique on flag information that makes our FTQEC protocols compatible with any fault-tolerant Clifford computation. (4) We use our optimization tools and perform numerical simulations on the hexagonal color codes [41] of distances 3, 5, 7, and 9. The results show that each of our tools can significantly reduce the logical error rates and increase the pseudothreshold for each code while preserving the code distance. For the hexagonal color code of distance 9, the pseudothreshold is improved by one order of magnitude, from $(1.34\pm 0.01)\times 10^{-4}$ to $(1.42\pm 0.12)\times 10^{-3}$, when all techniques are applied.
@@ III.1 Compact lookup table for minimum weight decoding and fault code
In this section, we discuss how to construct the fault set $\mathcal{F}_{t}$, verify its distinguishability, and construct the lookup table for error decoding. With our method, we can reduce the memory footprint requirement of the lookup table by 87.5% for self-orthogonal CSS codes compared to a lookup table designed for generic stabilizer codes. We also present the framework of fault codes that enables fast construction using streamlined Pauli-frame simulation represented as matrix algebra operations over $\mathrm{GF}(2)$.
@@ III.1.1 Reducing the memory footprint
Leveraging the structure of CSS codes, we can significantly improve the memory footprint. Assuming standard CSS decoding in which two separate lookup tables are used for $X$ and $Z$ decoding. Denote with $r_{X}$ and $r_{Z}$ the number of $X$- and $Z$-type stabilizer generators satisfying $r_{X}+r_{Z}=n-k$. The per entry cost decreases, as the entries only need to cater for $X$- or $Z$-type operators and syndromes. Each entry for the $X$- and $Z$-type syndromes will have $2r_{X}$ and $2r_{Z}$ bits respectively for the syndrome and the cumulative flag vector, and $n$ bits for the recovery operator. A self-orthogonal CSS code needs only one table futher decreasing the cost; see more detail on the total number of bits in Appendix A. Moreover, we can reduce the number of bits for the recovery operator to $k$ using the following two key ideas:
@@
    Given a syndrome $\vec{s}(E)$, the goal of decoding is to find a recovery operator $R$ such that $RE\in\mathcal{S}$, thus $R$ converts the error into the logical identity operation. For any possible Pauli error, we only have to store its logical class, a value that indicates how the error is related to a CRO with the same syndrome. This enables the map value to be only $2k$ bits of information in general and $k$ bits in case the code is a self-orthogonal CSS code. In this latter case and with $k=1$, the logical class is 0 if the multiplication of the Pauli operator and the CRO with the same syndrome is in the stabilizer group, otherwise the logical class is 1.
Altogether, for a self-orthogonal CSS code with $n\gg k$, the size of the table can be as small as 12.5% of the table if we viewed the code as a generic stabilizer code and stored the full recovery operators instead of the logical classes. For a CSS code that is not self-orthogonal, the gain is smaller but still significant. See Appendix A for detailed calculations around savings for the lookup table. Note that if the lookup table is used for proving distinguishability, all unique syndromes are required. However, in a realtime decoding architecture, the entries corresponding to significantly low probability fault combinations may be excluded, resulting in further reduction [47].
@@ III.1.2 Constructing the lookup table
Enumerating weight-1 faults— From here on, we will only consider a self-orthogonal CSS code, and denote its parity check matrix $H$. In order to list all possible single faults under the circuit-level depolarizing noise model, it is sufficient to consider all possible weight-1 faults within a single round of syndrome measurements. Each column of the fault check matrix $H_{\text{f}}$ describes for each possible weight-1 fault what its full syndrome and its logical class are. As the logical class of each fault depends on how its CRO is defined, we define the fault check matrix relative to the right inverse $H^{-1}$ of $H$ (for which $HH^{-1}=I_{(n-k)/2}$).
@@
Generalizing $H_{\text{f}}$ for a non-self-orthogonal CSS code is straightforward. In that case, the parity check matrices for $X$- and $Z$-type errors can be different, leading to different fault check matrices. Generalizing $H_{\text{f}}$ for a generic stabilizer code is more complicated but still doable, as all operators must be considered in the symplectic form. In that case, the number of rows for the logical class is $2k$. Also, instead of taking the inner product with $J_{i}$, whether each CRO commutes or anticommutes with each logical operator can be determined by the symplectic inner product between the symplectic bitstrings representing the CRO and the logical operator.
In the case that the code is a self-orthogonal CSS code, $n$ is odd, $k=1$, and logical $X$ and logical $Z$ operators are transversal, the fault check matrix is,
$H_{\text{f}}=\left(\begin{array}[]{c|c|c}H&0&\hbox{\multirowsetup$H_{\text{f,gate}}$}\\ 0&I_{(n-1)/2}&\\ J^{T}(I_{n}\oplus H^{-1}H)&0&\end{array}\right),$
@@
For errors of weight 1 on the data qubits, the operator $RE$ of each error can be represented by each column of $I_{n}\oplus H^{-1}H$. Since the logical class of $RE$ can be determined by its weight parity, the logical classes of this type of errors are the row of $L\equiv J^{T}(I_{n}\oplus H^{-1}H)$ where $J$ is the all-one column vector. That is for the Steane code, the part of $H_{\text{f}}$ corresponding to pure data qubit errors is
@@
$P_{i}=\begin{pmatrix}0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\\ 1&0&0&0&0&0\\ 0&0&0&1&0&0\\ 0&0&1&0&0&0\\ 0&0&0&0&0&1\\ 0&1&0&0&1&0\\ 0&0&0&0&0&0\\ 0&0&0&0&0&0\end{pmatrix}.$ (8)
@@
$PA=\begin{pmatrix}\Omega\\ \hline\cr\Phi\end{pmatrix}=\begin{pmatrix}\Omega_{1}&\Omega_{2}&\ldots&\Omega_{r}\\ \hline\cr\Phi_{1}&\Phi_{2}&\ldots&\Phi_{r}\end{pmatrix},$ (10)
@@
$L=J^{T}(\Omega\oplus\Theta).$ (13)
As a result, the part of $H_{\text{f}}$ corresponding to the gate faults is,
$H_{\text{f,gate}}=\begin{pmatrix}S\\ \hline\cr\Phi\\ \hline\cr L\end{pmatrix}=\begin{pmatrix}H\Omega\\ \hline\cr\Phi\\ \hline\cr J^{T}(I_{n}\oplus H^{-1}H)\Omega\end{pmatrix}.$ (14)
The relationship between the full syndrome, the data error, and the CRO of each fault is as follows: Suppose that the $i$-th column of $H_{\text{f}}$ (which represents a single fault on the $i$-th location) contains error syndrome $\vec{s}_{i}$, flag vector $\vec{f}_{i}$, and logical class $l_{i}$. The CRO of the fault is $H^{-1}\vec{s}_{i}$, while the data error of the fault is $l_{i}J\oplus H^{-1}\vec{s}_{i}$. That is, in case of a single fault, the actual recovery operator (ARO) we need to apply when finding the full syndrome $(\vec{s}_{i},\vec{f}_{i})$ is $l_{i}J\oplus H^{-1}\vec{s}_{i}$.
@@
When $t\geq 2$, we populate the cache with the logical classes of higher-weight fault combinations by simply combining all possible fault combinations of lower-weight fault combinations while keeping track of the weights of the fault combinations. We describe the $i$-th fault combination as a key-value pair $[(\vec{s}_{i},\vec{F_{i}}):(l_{i},w_{i})]$ where $(\vec{s}_{i},\vec{F_{i}})$ is the full syndrome, $l_{i}$ is the logical class, and $w_{i}$ is the weight of the fault combination. Combining the $i$-th and the $j$-th fault combinations gives $[(\vec{s}_{i}\oplus\vec{s}_{j},\vec{F_{i}}\oplus\vec{F_{j}}):(l_{i}\oplus l_{j},w_{i}+w_{j})]$. As we aim to check whether $\mathcal{F}_{t}$ is distinguishable, we fill up the cache by combining any pair of fault combinations that satisfy $w_{i}+w_{j}\leq t$. In case that the process gives the new key (the full syndrome) that already exists in the cache, we have a key conflict. This can be one of the following cases:
@@
If at the end we find that $\mathcal{F}_{t}$ is distinguishable, we can construct a lookup table of search radius $t$ from the cache as follows: for each key-value pair $[(\vec{s}_{i},\vec{F_{i}}):(l_{i},w_{i})]$ in the cache, we store a new key-value pair $[(\vec{s}_{i},\vec{F_{i}}):l_{i}J\oplus H^{-1}\vec{s}_{i}]$ in the lookup table (the weights are not necessary for decoding, though it might be useful for estimating the number of faults that causes the full syndrome). That is, $l_{i}J\oplus H^{-1}\vec{s}_{i}$ is the ARO for the full syndrome $(\vec{s}_{i},\vec{F_{i}})$. When performing error decoding, the ARO is applied if the full syndrome obtained from measurements is found on the lookup table; otherwise, the CRO ($H^{-1}\vec{s}_{i}$) is applied.
@@ III.1.3 The fault code
The perspective of the fault code can also be useful to extend a technique frequently used for error sampling (in qecsim [54], for example) to the circuit-level noise model beyond the code capacity noise model (memory errors only) and phenomenological noise model (both memory and measurement errors). Here, a randomly generated column vector of Hamming weight $w$ now represents faults on $w$ locations instead of errors on $w$ qubits. Suppose that the vector $\vec{v}$ represents the fault combination and $H_{\text{f}}\vec{v}$ gives the full syndrome $(\vec{s}(\vec{v}),\vec{F}(\vec{v}))$ and the logical class $l(\vec{v})$. In an error correction simulation, the decoder can predict the recovery operator $\vec{r}$ based on the full syndrome. We will find that the predicted recovery operator causes a logical error if and only if $l(\vec{v})$ and $l(\vec{r})$ differ.
@@ IV Optimization Tools for Time decoding
In this section, we develop several types of time decoders for flag FTQEC, building on the ideas of adaptive decoders for Shor-style error correction [38]. Different time decoders use different fault count estimation procedures. In Section IV.1, we describe a conventional way to perform repeated syndrome measurements for flag FTQEC in terms of difference vectors, which will be useful for the development in latter sections. In Section IV.2, we develop one-tailed and two-tailed adaptive time decoders which utilize flag information in the protocols. One-tailed adaptive decoder is applicable to a larger family of codes, while the two-tailed adaptive decoder is more optimized to self-orthogonal CSS codes but need to be used with an extended technique so that it becomes fully fault tolerant when applying to quantum computation. In Section IV.3, we develop two extended techniques that can further improve the performance of our adaptive time decoders for FTQEC, given that the code being used is a self-orthogonal CSS code.